OpenStudy (anonymous):

The body of a dump truck is raised to empty a load of dirt. How high is the 10-inch body from the frame when it is tipped up at 30 Degrees? Round to the nearest inch. A. 5 inches B. 7 inches C. 9 inches D. 11 inches

OpenStudy (anonymous):

This is dealing with Special Right Triangles Also this is my last question so take your time since I'm almost finished and have 50 minutes left in the test.

OpenStudy (anonymous):

Here's a pic for reference

OpenStudy (anonymous):

okaii so all you have to do is divide a number with another number that makes it to the 30 degrees then the second number you divided by is your answer

OpenStudy (anonymous):

Ok. So which numbers do I use? I've been through the directions and everything ad I still don't really understand it

OpenStudy (anonymous):

Which part of the truck is 10 inches?

OpenStudy (anonymous):

The bucket

OpenStudy (anonymous):

'10-inch body' could be referring to the hypotenuse or the horizontal leg. I would think it's the hypotenuse.

OpenStudy (anonymous):

Oh, ok, so you have the triangle and you want to find the missing side. You have the trig relation \[\sin30^\circ=\frac{10}{x}\] Solve for \(x\).

OpenStudy (anonymous):

So it's Sin 30 times 10?

OpenStudy (anonymous):

Not quite, you would multiply both sides by \(x\): \[x\sin30^\circ=\frac{10}{x}\cdot x\\ x\sin30^\circ=10\]

OpenStudy (anonymous):

Ok, gimme a sec

OpenStudy (anonymous):

Ok, so I got 5 for my answer, is that right?

OpenStudy (anonymous):

\[x=\frac{10}{\sin30^\circ}=\frac{10}{\frac{1}{2}}\not=5\]

OpenStudy (anonymous):

SO IT'S RIGHT :D

OpenStudy (anonymous):

No, it's not. Recheck your work.

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

\[\frac{10\cdot2}{\frac{1}{2}\cdot2}=\cdots\]

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

What's 10 times 2? What's 1/2 times 2?

OpenStudy (anonymous):

20 and 1

OpenStudy (anonymous):

Right, so \[\frac{10\cdot2}{\frac{1}{2}\cdot2}=\frac{20}{1}=20\]

OpenStudy (anonymous):

Ok, so what's the next step?

OpenStudy (anonymous):

Wait a second... I had the wrong equation at the start. It should have been \[\sin30^\circ=\frac{x}{10}\] Sorry about that... Your answer of 5 was correct.

OpenStudy (anonymous):

YAY!!! And it's ok XD Thanks for the help

OpenStudy (anonymous):

yw